A school of 300 students has a ratio of 3:2 girls to boys.What percentage are boys?
Answer by fcabanski(1391)

Call the number of any student x. According to the ratio there are 3x girls and 2x boys. Their number must equal 300:
3x + 2x = 300
5x=300
x=60
Boys is 2x = 120
Girls is 3x=180


% = part/whole = boys/students = 120/300 = .4 = 40%

If you need help understanding math so you can solve these problems yourself, then one on one online tutoring is the answer ($30/hr). If you need faster solutions with guaranteed detailed answers, then go with personal problem solving ($3.50-$5.50 per problem). Contact me at fcabanski@hotmail.com